OpenStudy (anonymous):

Max and jack each throw a die. find the probability that: (a) the number thrown by max is greater than the number thrown by jack.

OpenStudy (anonymous):

there are many possible cases

OpenStudy (anonymous):

let jack thrown 1, then max can throw 2,3,4,5,6

OpenStudy (anonymous):

i understand that but how do you work out the probability for it?

OpenStudy (anonymous):

i tried multiplying the possibilities but the answer wasn't right

OpenStudy (anonymous):

the answer should be = 5/36+8/36+9/36+8/36+5/36 =35/36

OpenStudy (anonymous):

thats not the answer. its 5/12

OpenStudy (anonymous):

thats not the answer. its 5/12

OpenStudy (anonymous):

thats not the answer. its 5/12

OpenStudy (anonymous):

count

OpenStudy (anonymous):

it is only the statement of this question that makes it confusing. suppose you just said "toss a red and a green die. what is the probability that the green one is larger?"

OpenStudy (across):

Jack can throw one of these: 1, 2, 3, 4, 5, 6 For each of Jack's throws, there's a probability that Max will throw a larger number associated to it, as follows: 1 - 5/6 2 - 2/3 3 - 1/2 4 - 1/3 5 - 1/6 Add those probabilities and you'll obtain: 1/2

OpenStudy (anonymous):

1/2 isn't the answer though

OpenStudy (anonymous):

roll two dice, 36 possible outcomes. the ones where the second one is bigger are \ (1,2).(1,3).(1.4),(1,5),(1,6) (2,3),(2,4),(2,5) (2,6) (3,4),(3,5),(3,6) (4,5),(4,6) (5,6)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

when i count them i get 15, and \[\frac{15}{36}=\frac{5}{12}\] is the answer i get

OpenStudy (anonymous):

thats correct thank you:)

OpenStudy (anonymous):

remember when you use 2 dice you do not need any fancy formulas. only 36 elements, just count the ones you want. you can use it to CHECK formula, but you can always count!

OpenStudy (anonymous):

ok got it:) thank you

OpenStudy (anonymous):

yw